A solid body starts roting about a stationary axis with an angular velocity=t/50,t is in sec
how soon after the beginning of rotation will the total acceleration vector of an arbitrary point of the body form an an angle 60degrees with its velocity vector?

Grade:

1 Answers

vikas askiitian expert
13 years ago

v=t/50  ( given)

dv/dt = at(tangential)=1/50

ar(centripital)=v^2/r = (t^2/2500r)

total accleration = sqrt ( (at)^2 + (ar)^2 )

 total accleration is making an angle 60 with velocity vector ......velocity is parallel to tangential accleration so total accleration makes angle of 60 with tangential accleration...

a(total accleration)cos60 = a(tangential)             

t= sqrt(3r) sec        ( here r is the radius of circular path)